Reading List for the PhD Qualify Examination

School of Computer and Information Science

Student: Gang Cheng

Dissertation Advisor: Prof. Geoffrey C. Fox

Committee Members: Prof. Nancy McCracken, Prof. Salim Hariri

References

[1] George F. Coulouris, Jean Dollimore, Distributed Systems: Concepts and Design, Addison-Wesley, 1988.

[2] G. Fox, M. Johnson, G. Lyzenga, S. Otto, J. Salmon, D. Walker, Solving Problems on Concurrent Processors, Vol. 1, Prentice Hall, Englewood Cliffs, N.J., 1988.

[3] A. Ananda and B. Srinivasan, IEEE Tutorial: Distributed Computing Systems: Concepts and Structures, IEEE Computer Society Press, 1991.

[4] System Software and Tools for High Performance Computing Environments, Final report of the Workshop on System Software and Tools for High Performance Computing Environments, Pasadena, California, April 14-16, 1992, 225 pages.

[5] R. M. Karp, R. E. Miller, Properties of a Model for Parallel Computations: Determinacy, Termination, Queueing, SIAM J. Appl. Math., Vol. 14, No. 6, Nov. 1966.

[6] Paul A. Suhler, J. Biswas, K. M. Korner, J. Browne, TDFL: A Task-Level Dataflow Language, Journal of Parallel and Distributed Computing 9, 103-115(1990).

[7] D. E. Culler, G. K. Maa, Assessing the Benefits of Fine-grain Parallelism in Dataflow Programs, The International Journal of Supercomputer Applications, Vol. 2, No. 3, 1988, pp. 10-36.

[8] G. Fox, S. Hiranadani, K. Kennedy, C. Koelbel, U. Kremer, C-W Tseng, and M-Y Wu, Fortran D Language Specification, Syracuse Center for Computational Science-42c, Rice COMP TR90-141, 37 pps, 1991.

[9] I. Foster and K. M. Chandy, Fortran M: A Language for Modular Parallel Programming, Preprint MCS-P237-0992, Mathematics and Computer Science Division, Argonne National Laboratory, Argonne, Ill., 1992.

[10] K. Mani Chandy, C. Kesselman, Compositional C++: Compositional Parallel Programing, Technical Report, California Institute of Technology, 1992.

[11] V. Sunderam, PVM: A Framework for Parallel Distributed Computing, Concurrency: practice and experience, 2(4), Dec. 1990.

[12] A. Beguelin, J. Dongarra, G. A. Geist, R. Manckek, V. S. Sunderam, Graphical Development Tools for Network-Based Concurrent Supercomputing, Proc. of Supercomputing ’91, IEEE, pp. 435-445, 1991.

[13] L. L. Smarr, Metacomputing, Communication of the ACM, Vol. 35, No.6, June, 1992.

[14] S. Hariri and G. Fox, Trends Toward Virtual Computing Environment, Technical Report, NPAC, Syracuse University, March, 1993.

[15] Craig Upson, Thomas Faulhaber, Jr., David Kamins, David Laidlaw, David Schlegel,

2 Reading List

Jeffrey Vroom, Robert Gurwitz and Andries van Dam, The Application Visulization System: A Computational Environment for Scientific Visualization, IEEE Computer Graphics and Applications, July, 1989.

[16] G. Oberbrunner, Parallel Networking and Visualization on the Connection Machine CM-5, the Symposium on High Performance Distributed Computing HPDC-1, September, 1992, pp. 78-84, Syracuse, NY.

[17] C. King, W. Chou and M. Ni, Pipelined Data-Parallel Algorithms: Part I - Concept and Modeling, IEEE Transactions on Parallel and Distributed Systems, Vol. 1, No. 4, October